OpenStudy (anonymous):

what are the steps to calculate sin20 /cos70

OpenStudy (anonymous):

without a calculator, you mean?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

gve me the steps?

OpenStudy (anonymous):

Oh, lol, I thought it would be some boring addition/subtraction formulae, but I forgot a key fact: sin(x) = cos(90 - x) for all x.

OpenStudy (anonymous):

Glad, too, as 20 and 70 aren't the easier to calculate directly (but it can be done without THAT much hastle)

OpenStudy (anonymous):

sin(20)=cos(90-20)=cos(70). it's pretty obvious now, isn't it?

OpenStudy (anonymous):

It was obvious before that, tbh. I could just sense it.

OpenStudy (anonymous):

LOL!! I am sure it was obvious for you.. I was talking to leonasmart!

OpenStudy (anonymous):

You're looking for a fight, aren't you Michael?

OpenStudy (anonymous):

No.. :(

OpenStudy (anonymous):

lol...thx

OpenStudy (anonymous):

so wats the answer

OpenStudy (anonymous):

Michael, What about you post a real challenging problem if you have one?

OpenStudy (anonymous):

Do you think you could answer it?

OpenStudy (anonymous):

I posted one of my 'fun' problems yesterday, btw. No one got it, so I had to give the solution.

OpenStudy (anonymous):

I used the formula that Newton wrote above to show that sin(20) is the same as cos(70), therefore, \[{\sin(20) \over \cos(70)}={\cos(70) \over \cos(70)}=1\]

OpenStudy (anonymous):

What was the problem?

OpenStudy (anonymous):

ok...thx again

OpenStudy (anonymous):

Oh, nothing too serious, just some simultaneous equations - is it possible to link to old topics?

OpenStudy (anonymous):

You're welcome leonasmart!

OpenStudy (anonymous):

Yeah I guess.

OpenStudy (anonymous):

i am going to try do one on my own now and ask u if it is correct

OpenStudy (anonymous):

btw leonas if you want proof of that sin(90-x) thing consider a RA triangle.

OpenStudy (anonymous):

OK.. We are around.

OpenStudy (anonymous):

ok.....go on INewton

OpenStudy (anonymous):

Congrats for the new title Newton<< (like you would care :P)

OpenStudy (anonymous):

:D I do care

OpenStudy (anonymous):

OK WHAT about sin 30/cos 60

OpenStudy (anonymous):

You do?

OpenStudy (anonymous):

the same as the old one.

OpenStudy (anonymous):

my question (but answer is there so you can cheat :( ) http://openstudy.com/updates/4dac7055d6938b0b7bb9a74d#/updates/4dac7055d6938b0b7bb9a74d hope links work here. Note, I can ask some MUCH harder ones if you want.

OpenStudy (anonymous):

OK.. I won't cheat :)

OpenStudy (anonymous):

ok thx

OpenStudy (anonymous):

if I got the answer, I would post it here..K?

OpenStudy (anonymous):

Ok then.

OpenStudy (2bornot2b):

Can you please help me with my problem at http://openstudy.com/groups/mathematics#/updates/4dadd133d6938b0b6b21a94d

OpenStudy (anonymous):

about the hint you gave, what if x+y is 0?

OpenStudy (anonymous):

It makes no difference (but if they were the rest wouldn't fit anyway, so you can forget that possibility, as they cannot satisfy it). Another hint is to (if you realise why multiplying is helpful), to apply the same thing to the third one.

OpenStudy (anonymous):

I already did, and I almost got the values of x and y, but the numbers don't look right to me. They are a little bit complicated. So, I am doing it again.

OpenStudy (anonymous):

No more hints Please!

OpenStudy (anonymous):

There are radicals in the answer :/ Sorry.

OpenStudy (anonymous):

Oh really?

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

Yes. The answers aren't nice, but the idea is good (most questions of the same type (same source) are much less dirty maths though, but it is probably one of the easiest to get into so I wanted to test the water with it). It's from a non calculator exam, though, so nothing TOO nasty.

OpenStudy (anonymous):

So far, I have the two equations, which can find x and y: x+y=48/35 (1) and xy=9/35.

OpenStudy (anonymous):

Yes, I'm not a big fan of irrational numbers either.

OpenStudy (anonymous):

Hmm, your equations are slightly inaccurate.

OpenStudy (anonymous):

But you are meant to obtain values for (x+y) and xy first, so your method is obviously right.

OpenStudy (anonymous):

I know my method is right :P .. I want my answer to be right!! just gimme a minute to check my calculations.

OpenStudy (anonymous):

hmm ok. While we're chatting, can I ask you a question: Just HOW good at Maths is Lokisan? If you can answer that.

OpenStudy (anonymous):

Well I don't know that much about him.. But, from what I have seen so far, I can say he's pretty good. You can't tell if he's special since there are not so many challenging problems here.. I think you would agree with that.

OpenStudy (anonymous):

I do agree with that, actually (mainly the reason I asked, as I have no way to find out). The only problem is no one on the site really wants challenging problems.

OpenStudy (anonymous):

What about we two make a change?

OpenStudy (anonymous):

Haha, I may post another 'fun' question later (but this time with nicer maths and not horrid answers). But unfortunately I don't think many others care :(

OpenStudy (anonymous):

You didn't tell me what you think about lucas?

OpenStudy (anonymous):

Oh, I agree, he's obviously good. I pretty much think the same about not being able to judge on here though (he 'knows' higher level maths than me (I am much younger I am almost sure) but unsure exactly how good he is).

OpenStudy (anonymous):

May I ask how old you are?

OpenStudy (anonymous):

19. Yourself?

OpenStudy (anonymous):

21

OpenStudy (anonymous):

xy=3/35.. this can't be wrong.

OpenStudy (anonymous):

It's right

OpenStudy (anonymous):

x+y=30/35?

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

35xy=3 and 35x+35y=30 implies x(30-35x)=3

OpenStudy (anonymous):

Am I allowed to use software to solve the quadratic equation? :(

OpenStudy (anonymous):

yeah, sure (it's boring from there)

OpenStudy (anonymous):

Really? Did I already get all the fun? :(

OpenStudy (anonymous):

Obviously I did. :(

OpenStudy (anonymous):

:( I'll keep a lookout for a nicer one

OpenStudy (anonymous):

You know I will get two values for x and y, substitute them in two of the four equations to get a and b.

OpenStudy (anonymous):

Yeah (x/y and a/b are interchangeable now, except their signs compliment each other)

OpenStudy (anonymous):

I actually have a nice question I just found, but I won't have been able to type it up for ~ 10/15 minutes

OpenStudy (anonymous):

This is much nicer (it's more if you see the trick it's a very quick, nice problem, rather than grinding even after spotting how to do it)

OpenStudy (anonymous):

Waiting for it :)

OpenStudy (anonymous):

I wonder why nobody solved this problem. It's not that difficult.

OpenStudy (anonymous):

Exactly. I guess it's longer than most though. And most people here do problems to help people more than for just doing them.

OpenStudy (anonymous):

Yeah. I usually don't solve problems that take too much time.

OpenStudy (anonymous):

What major are you taking? ( assuming you already go to college )

OpenStudy (anonymous):

Well University, but in England, so no 'major' - it's a slightly different system - just Maths.

OpenStudy (anonymous):

:O you don't have majors in England.. So you're going for a degree in Mathematics.

OpenStudy (anonymous):

Indeed. Ugh I have the question but there is a problem in the LaTeX. Trying to find it.

OpenStudy (anonymous):

It's 10:26 PM at where I am, and I didn't get enough sleep last night. So, if it would take me too long, I might delay solving it to tomorrow morning.

OpenStudy (anonymous):

Let \[I = \int^a_0 \frac{f(x)}{f(x)+f(a-x)} \mathbb{d}x\] Use a substitution to show that \[I = \int^a_0 = \frac{f(a-x)}{f(x)+f(a-x)}\mathbb{d}x \] and hence evaluate I in terms of a. Use this result to evaluate the integrals \[\int^1_0 \frac{ln(x+1)}{ln(2+x-x^2)}\mathbb{d}x\ \] and \[\int^{\frac{\pi}{2}}_0 \frac{\sin x}{\sin (x + \pi/4)} \mathbb{d}x \] Evaluate \[\int^2_{\frac{1}{2}} \frac{\sin x}{x(\sin x + \sin 1/x)} \]

OpenStudy (anonymous):

OK. I think that's right but if I see a typo I'll let you know

OpenStudy (anonymous):

There is a typo. at least one.

OpenStudy (anonymous):

Where? The last integral need a dx, but apart from that..?

OpenStudy (anonymous):

2nd one

OpenStudy (anonymous):

Ah, an extra = sign :/ oops

OpenStudy (anonymous):

Are you sure about the second integral? I think something is wrong.

OpenStudy (anonymous):

For which part?

OpenStudy (anonymous):

f(a-x) / [(f(x) + f(a-x)] ?

OpenStudy (anonymous):

is the substitution a-x?

OpenStudy (anonymous):

Yes, u = a-x (or whatever you call the variable) Note u and x are just dummy variables, so if you get the same thing in u as the one in x it is the same integral.

OpenStudy (anonymous):

Oh Yeah.. That's what I was stuck at :(

OpenStudy (anonymous):

whose good i writing essays

OpenStudy (anonymous):

I am, but not for you, Mary.

OpenStudy (anonymous):

Hint, though Mary: It's ' Who's' (short for who is) not whose.

OpenStudy (anonymous):

can you help me write it